OpenStudy (anonymous):

Perform the operation as indicated and express answers in simplified form. The square root of 27x to the power of 5

zepdrix (zepdrix):

\[\sqrt{27x^5}\]\[(\sqrt{27x})^5\] which way is it written? :D

OpenStudy (anonymous):

with the five inside the brackets.

OpenStudy (anonymous):

like the top one.

OpenStudy (jiteshmeghwal9):

Power of 5=\(x^5\) \((\sqrt{27x})^5\)

OpenStudy (ajprincess):

\(\sqrt{27x^5}=\sqrt{3^2*3*x^2*x^2*x}\) \(=3x^2\sqrt{3*x}\) \(=3x^2\sqrt{3x}\)